OpenStudy (solomonzelman):

Dirichlet's Test, 2 examples.

OpenStudy (solomonzelman):

Series: \(\color{#000000 }{ \displaystyle \sum_{t=s}^\infty \frac{\cos t}{\ln t} }\) Contains: (1) General Theory of Dirichlet's Test (2) Derivation of partial sum of the geometric series Purpose: (!) Proof that the series converges, by Dirichlet's Test. (This is the final version, with more precise wording)

OpenStudy (solomonzelman):

Series: \(\color{#000000 }{ \displaystyle \sum_{t=s}^\infty t^{-0.0001}\sin t }\) Purpose: (!) Proof that the series converges, by Dirichlet's Test. (I don't repeat the theory part of the geometric series and Dirichlet's Test in document #2, and additionally I am informing you that THIS is the final version (of doc 2), with more precise wording)